Ahmed wants to pay back his debt to the bank. He has already paid back 60% of the initial debt. His family offered to help by paying 33% of the initial debt, so now he only needs an additional $105 in order to pay off the debt. What was the initial amount of the debt Chris owed to the bank? *
Answer by jim_thompson5910(35256)

"He has already paid back 60% of the initial debt. His family offered to help by paying 33% of the initial debt"

So he has 60% + 33% = 93% of the debt taken care of leaving 100% - 93% = 7% remaining

That 7% is represented by $105 since he needs this amount to fully pay off the debt.

Let x be the amount of total debt initially owed.
7% of x is the same as saying 0.07*x since 7% = 7/100 = 0.07
the "of" means "multiply"

Set 0.07x equal to 105 and solve for x
0.07x = 105
0.07x/0.07 = 105/0.07 <<-- divide both sides by 0.07
x = 1500

------------------------------------------
------------------------------------------
Answer:
The initial amount of the debt was $1500